# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ Abstract
This article provides a tutorial on over-the-air electromagnetic signal processing (ESP) for next-generation wireless networks, addressing the limitations of digital processing to enhance the efficiency and sustainability of future 6th Generation (6G) systems. It explores the integration of electromagnetism and signal processing (SP) highlighting how their convergence can drive innovations for 6G technologies. Key topics include electromagnetic (EM) wave-based processing, the application of metamaterials and advanced antennas to optimize EM field manipulation with a reduced number of radiofrequency chains, and their applications in holographic multiple-input multiple-output systems. By showcasing enabling technologies and use cases, the article demonstrates how wave-based processing can minimize energy consumption, complexity, and latency, offering an effective framework for more sustainable and efficient wireless systems. This article aims to assist researchers and professionals in integrating advanced EM technologies with conventional SP methods.
@@ I Introduction
After the publication of Shannon’s landmark paper in 1948, the digital revolution rapidly accelerated, shifting information processing and communication from the analog realm to the digital domain. The exponential growth in digital computing capabilities has pushed the boundary between analog and digital processing in wireless systems closer to the devices’ antennas. This transition is mainly driven by software-defined radio implementations, which have shifted most analog processing to radio-frequency (RF) circuits, primarily utilizing guided wave electromagnetic (EM) structures. Except for optical systems, where analog computing has been established for several years and remains an active research area [1], most signal processing in wireless systems occurs in the digital domain at baseband. In these systems, RF circuits are primarily responsible for basic operations such as up-down conversion, phase shifts, and amplification. However, this approach could be questioned when addressing the design of next-generation wireless systems, such as 6th Generation (6G) networks, which are expected to impose extremely stringent requirements to offer enhanced performance in terms of capacity, reliability, reduced latency, and new functionalities like integrated sensing and communications (ISAC) [2, 3]. So far, these needs have been addressed using various solutions, such as small cells, millimiter wave (mmWave) communications, and massive multiple-input multiple-output (MIMO) systems, which have demonstrated their ability to improve network capacity and provide ubiquitous connectivity [4].
Nonetheless, these well-established technologies and design practices no longer appear sufficient to achieve the performance levels required for future wireless networks. In fact, the evolution towards 6G is driving the exploration of the fundamental limits of wireless communication and innovative design paradigms to meet the stringent requirements posed by next-generation networks [4, 5]. For instance, the use of higher frequency bands, e.g., terahertz (THz), and electrically large antenna arrays, e.g., based on metasurfaces and extremely large-scale multiple-input multiple-output (XL-MIMO) configurations, are key enablers of this evolution [6]. Interestingly, their joint utilization enables the system’s operation in the radiative near-field region of the antennas, thus leading to unprecedented levels of communication and sensing performance, flexibility, and resolution [7, 8]. However, we have reached the point where further advancing the capabilities of already established technologies and incorporating new ones is hindered by issues related to practical implementation, hardware cost, and complexity, as well as excessive processing time and power consumption. These issues often result in the so-called “digital bottleneck”, posing a substantial challenge to the feasibility, scalability, and sustainability of future wireless networks.
Therefore, it is increasingly evident that it is necessary to reverse the current trend and conceive solutions where at least part of the signal processing occurs in the analog domain at RF. For instance, hybrid digital-analog solutions in MIMO systems have been extensively explored to mitigate digital processing burdens, though often at the cost of reduced flexibility [4]. Nevertheless, current hybrid solutions are not sufficient and radically new physical layer technologies, featuring a more balanced integration between digital and analog RF processing, have to be found.
@@
A promising approach toward this goal is to transfer part of the signals’ processing directly to the EM level [9, 10, 11, 12, 13, 14]. This approach is known as over-the-air electromagnetic signal processing (ESP), also referred to as electromagnetic signal and information theory (ESIT) or electromagnetic information theory (EIT). Specifically, ESP can be defined as an interdisciplinary scientific field focused on exploiting the synergies between innovative EM technologies, signal processing (SP) algorithms, and information theory. This term refers to the study and development of novel technological solutions where an EM wave generated by one or more radiating sources is perturbed during propagation, i.e., over the air, by reconfigurable EM structures. This approach aims to achieve specific SP functions with lower energy consumption, complexity, and latency (the processing takes place at the speed of light) compared to their digital counterparts. In addition, ESP can be considered as part of the more general holographic communications framework, which is envisioned as a holistic way to manipulate the EM field generated or sensed by the antennas with unprecedented flexibility [8]. This paradigm involves designing reconfigurable radio propagation environments [15] that can arbitrarily control signal transmission, manipulation, and reflection through the massive deployment of peculiar EM devices based on metamaterials. In these devices, the reflection, refraction, or scattering properties of a set of reconfigurable scatterers composing their structure are optimized to achieve specific processing and radiation characteristics.
@@
Admittedly, many of the studies on the design of innovative EM structures and processing algorithms are still in their infancy, and significant progress is needed in terms of practical implementations and concrete applications. The seamless integration of electromagnetism and SP offers a transformative opportunity for advancing next-generation wireless networks, particularly through the emergence of ESP. As we progress towards 6G systems, the SP community encounters new challenges that require a reassessment of traditional methods and a shift towards innovative approaches.
@@
The primary objective of this article is to present in a unified way the fundamental limits and the new opportunities in processing EM waves using physically consistent models. While traditional SP techniques are effective in the digital realm, they often struggle with the complexities and constraints posed by EM environments and emerging EM technologies. This discrepancy underscores the necessity for specialized SP algorithms that align with the physical properties and constraints of EM waves. This article seeks to shed light on this issue and wants to provides a survey of recent advancements in ESP. It emphasizes the importance of bridging EM theory with SP and advocates for a unified strategy to enhance the sustainability of wireless communication technologies. Targeted at readers with expertise in SP and/or wireless communications, the article begins with a brief historical overview before revisiting fundamental EM principles needed to establish a general framework for exploring the fundamental limits and practical algorithms of ESP. Key technologies will be presented as specific cases, supported by numerical examples that underscore the need for electromagnetically consistent optimization algorithms. By addressing the limitations, opportunities, and challenges of this integrated ESP approach, we aim to inspire further research and provide valuable insights. We hope this work serves as a useful resource for researchers and practitioners, highlighting the essential role of integrating these disciplines in driving innovative solutions and applications for future 6G networks.
Figure 1: EM signal processing performed in the reactive near field (top) or in the radiative near field (or far-field) (bottom) regions of the sources using a reconfigurable scattering device.
@@
In a general setting, ESP refers to the capability of processing EM waves over the air by introducing one or more reconfigurable scattering devices between the source and receiver spaces, denoted as $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$, respectively (see Fig. 1). Typically, these reconfigurable scattering devices are linear and passive structures thus reducing their power consumption close to zero. Let ${\cal{S}}$ denote the space occupied by the scattering device. The configuration of the device can be modified using a set of parameters, $\bm{\theta}$, which vary depending on the specific implementation technology. The scenario shown in Fig. 1 can encompass various ESP technologies and geometric configurations. Specifically, ESP solutions can generally be categorized into two types:
@@
  • Reactive region configuration: The scattering device is located in the reactive region of the source space so that inductive coupling takes place between them (see Fig. 1, top). Its role includes enabling large antenna arrays while minimizing the number of RF chains, reducing power consumption, and lowering latency compared to traditional solutions. Additionally, it can support the implementation of complex functionalities in the analog domain. The dual scheme where the scattering device is located close to the receiver space is also possible.
@@
  • Radiative near and far field region configuration: The reconfigurable scattering device is positioned in the radiative region of the source and receiver spaces and it is used to control the propagation environment. This approach offers a cost-effective alternative to deploying additional access points (see Fig. 1, bottom). In some configurations, the source and receiver spaces might match (see Sec.V-C).
Regardless of the category above or specific technology, the core ESP problem can be formulated as follows. Given a set of sources located in the source space $\mathcal{S}_{\text{T}}$, determine the configuration $\bm{\theta}$ of the scattering device that generates the desired EM field at the receiver space $\mathcal{S}_{\text{R}}$, while satisfying the constraints imposed by EM laws and the selected technology. A further step involves making the set of parameters $\bm{\theta}$ dependent on data generated within the scattering device itself (e.g., by a sensor in an Internet of things (IoT) application). This allows the information to be directly embedded into the reflected signal, removing the need for additional RF signal generation. Consequently, this approach can significantly reduce energy consumption at the sensor side [16, 17, 18].
@@
Notably, the concept of reconfigurable scatterers has a long and well-established history in the field of antenna theory. A pivotal milestone and precursor to contemporary solutions is R. Harrington’s theory of loaded scatterers, which is closely associated with characteristic mode decomposition. This theory involves configuring scatterers such that the principal radiating mode of the EM structure, when illuminated by the source wave, aligns with the desired response [19]. This theory was later extended, also by R. Harrington, to reactively controlled arrays. In this context, passive scattering elements are positioned near a radiating antenna to exploit mutual coupling within the reactive near-field region, thereby modifying the antenna’s radiation characteristics [20].
In recent years, these concepts have been revisited and advanced thanks to technological and material innovations, enabling the development of new EM structures that address the increasing demand for enhanced performance, flexibility, and cost efficiency in future wireless systems. A notable example is given by metasurfaces, which are bi-dimensional (2D) structures, either homogeneous or inhomogeneous, composed of interacting sub-wavelength-sized cells called meta-atoms. These structures may provide control of the reflected and transmitted EM field [21, 22]. Both metasurfaces and more conventional solutions, such as backscattering antenna arrays, have facilitated the development of reconfigurable intelligent surfaces, which have garnered significant interest within the scientific community. RISs are primarily used to manipulate the propagation environment and are considered a cornerstone of the emerging smart radio environments paradigm, whose aim is to integrate the propagation environment into the system’s design and optimization loop [23, 24]. Specifically, RISs belong to the second category, shown in Fig. 1 (bottom), where they are typically employed to control the propagation environment [24, 25].
@@
Within the first category shown in Fig. 1 (top), there is a wide variety of promising technologies. These include XL-MIMO arrays (also known as extremely large antenna arrays) [26], large intelligent surfaces or holographic multiple-input multiple-output (HMIMO) surfaces [5, 27], dynamic metasurface antennas [7], fluid and reconfigurable antennas [28], and the latest developments in reactively controlled arrays, such as electrically steerable passive array radiators and dynamic scattering arrays [29, 30]. Significant advancements have concurrently been made in the optical community with the development of all-optical diffractive deep neural networks [31], and a similar concept has recently been applied to RF systems [32], inspiring the development of the stacked intelligent metasurface (SIM) technology, even though this technology is currently limited to linear processing capabilities [33]. Additionally, the introduction of novel volumetric metamaterials has enabled three-dimensional (3D) EM processing functions, such as first-derivative computation [34].
As discussed in the following sections, a key feature of novel EM structures from an SP perspective is their ability to perform complex operations at the EM level. Some of the processing functions that can be achieved include over-the-air computation of the singular value decomposition (SVD) of the MIMO channel matrix, spatial derivatives, MIMO EM precoding, and direction-of-arrival (DoA) signal estimation.

### v3

@@ Abstract
This article provides a tutorial on over-the-air electromagnetic signal processing (ESP) for next-generation wireless networks, addressing the limitations of digital processing to enhance the efficiency and sustainability of future 6th Generation (6G) systems. It explores the integration of electromagnetism and signal processing (SP) under a unified framework by highlighting how their convergence can drive innovations for 6G technologies. Key topics include electromagnetic (EM) wave-based processing, the application of metamaterials and advanced antennas to optimize EM field manipulation with a reduced number of radiofrequency chains, and their applications in holographic multiple-input multiple-output systems. By showcasing enabling technologies and use cases, the article illustrates how wave-based processing can minimize energy consumption, complexity, and latency, offering an effective framework for more sustainable and efficient wireless systems. This article aims to assist researchers and professionals in integrating advanced EM technologies with conventional SP methods.
@@ I Introduction
After the publication of Shannon’s landmark paper in 1948, the digital revolution rapidly accelerated, shifting information processing and communication from the analog realm to the digital domain. Since then, the exponential growth in digital computing capabilities has pushed the boundary between analog and digital processing in wireless systems increasingly closer to the devices’ antenna. This evolution is mainly driven by software-defined radio implementations, where analog processing is relegated to intermediate frequency (IF) and radio-frequency (RF) circuits, mainly for basic operations such as filtering, amplification, phase shifting, and frequency conversion. In fact, except for optical systems—where analog computing has been established for several years and remains an active research area [1]—most signal processing in current wireless systems takes place in the digital domain at baseband.
This approach, however, could be questioned when addressing the design of next-generation wireless systems, such as 6th Generation (6G) networks, which are expected to impose extremely stringent requirements to deliver enhanced performance in terms of capacity, reliability, latency, and support for new functionalities such as integrated sensing and communications (ISAC)[2, 3]. An analysis of the evolution of mobile radio technologies over recent decades reveals that the growing demands of each new generation—from 1G to 5G—have been addressed through the adoption of advanced solutions such as small cell deployments, millimeter wave (mmWave) communications, and massive multiple-input multiple-output (MIMO) systems. These innovations have effectively boosted network capacity, enhanced performance, and supported ubiquitous connectivity [4]. Looking ahead to the next generation, significant research efforts are focused on the use of higher frequency bands—such as terahertz (THz)—and electrically large antenna arrays, including metasurfaces and extremely large-scale multiple-input multiple-output (XL-MIMO) systems [5]. Notably, their combined deployment enables operation in the radiative near-field region of the antennas, unlocking unprecedented levels of communication and sensing performance, flexibility, and spatial resolution [6, 7].
However, we have reached the point where further advancing the capabilities of already established technologies and incorporating new ones is hindered by issues related to practical implementation, hardware cost, and complexity, as well as excessive processing time and power consumption. These issues often result in the so-called “digital bottleneck”, posing a substantial challenge to the feasibility, scalability, and sustainability of future wireless networks.
Increasingly, it is becoming clear that reversing the prevailing trend toward fully digital processing is necessary, prompting the design of solutions in which at least part of the signal processing is performed in the analog domain at RF. For instance, hybrid digital-analog solutions in MIMO systems have been extensively explored to mitigate digital processing burdens, though often at the cost of reduced flexibility [4]. Nevertheless, hybrid XL-MIMO systems, even with analog precoding and low-resolution data converters, entail substantial power consumption. Moreover, the proliferation of antennas and RF components further complicates hardware design [8]. Similarly, the integration of ISAC into future mobile networks, widely regarded as a key technological breakthrough, presents substantial challenges, particularly due to the requirement for transceivers to process signals with very high dynamic ranges, e.g., from strong communication signals to weak radar echoes. This necessitates high-resolution analog-to-digital converters, which significantly increases power consumption in both fully digital and hybrid architectures, potentially making ADCs and digital processing units the primary sources of energy consumption, surpassing that of the analog RF front-end [9]. Although low-resolution ADCs, including extreme one-bit quantization, can offer competitive performance in some applications when paired with advanced signal processing, their practical viability depends on factors such as computational complexity, processing unit cost, and potential strict constraints on responsiveness [3].
@@
It is precisely the growing awareness of the limitations of current technologies, coupled with the pressing need to reduce their energy impact, that is fueling the exploration of innovative design paradigms. A promising approach toward this goal is to transfer part of the signals’ processing directly to the electromagnetic (EM) level [10, 11, 12, 13, 14]. This approach is known as over-the-air electromagnetic signal processing (ESP), also referred to as electromagnetic signal and information theory (ESIT) or electromagnetic information theory (EIT). Specifically, ESP can be defined as an interdisciplinary scientific field focused on exploiting the synergies between innovative EM technologies, signal processing (SP) algorithms, and information theory. This term refers to the study and development of novel technological solutions where an EM wave, generated by one or more radiating sources, is perturbed during its propagation, i.e., over-the-air, by reconfigurable, typically passive, EM structures. This approach aims to perform specific SP functions with reduced energy consumption, complexity, and latency—since processing occurs at the speed of light—compared to their digital counterparts. In fact, ESP can be viewed as part of the broader holographic communications paradigm, which envisions a holistic method for manipulating the EM field generated or sensed by antennas with unprecedented flexibility [7]. This paradigm involves designing reconfigurable radio propagation environments [15] that can arbitrarily control signal transmission, manipulation, and reflection through the massive deployment of peculiar EM devices based on metamaterials. In these devices, the reflection, refraction, or scattering properties of a set of reconfigurable scatterers composing their structure are optimized to achieve specific processing and radiation characteristics.
@@
Admittedly, many studies on the design of innovative EM structures and processing algorithms are still in their infancy, and significant progress is needed in terms of practical implementations and concrete applications. Nonetheless, the seamless integration of electromagnetism and SP offers a transformative opportunity for advancing next-generation wireless networks, particularly through the emergence of ESP.
@@
The primary objective of this article is to present, in a unified way, the fundamental limits and new opportunities in processing EM waves using physically consistent models, establishing a bridge between typically fragmented frameworks such as EM, antenna, circuit, and SP theories. While traditional SP techniques are effective in the digital realm, they often struggle with the complexities and constraints posed by EM environments and emerging EM technologies. This discrepancy underscores the necessity for specialized SP algorithms that align with the physical properties and constraints of EM waves. This article seeks to shed light on this issue and provide a survey of recent advancements in ESP. It emphasizes the importance of bridging EM theory with SP and advocates for a unified strategy to enhance the sustainability of wireless communication technologies.
Targeted at readers with expertise in SP and/or wireless communications, the article begins with a brief historical overview before revisiting fundamental EM principles needed to establish a general framework for exploring the fundamental limits and practical algorithms of ESP. Key technologies will be presented as specific cases, supported by numerical examples that underscore the need for electromagnetically consistent optimization algorithms.
By addressing the limitations, opportunities, and challenges of this integrated ESP approach, we aim to inspire further research and provide valuable insights. We hope this work serves as a useful resource for researchers and practitioners, highlighting the essential role of integrating these disciplines in driving innovative solutions and applications for future 6G networks.
@@
In a general setting, ESP refers to the capability of processing EM waves over the air by introducing one (or many) reconfigurable scattering devices between the source and receiver spaces, denoted as $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$, respectively (see Fig. 1). The source space might represent a transmitter equipped with multiple active antennas, while the receiver space may correspond to one or more receivers.
Figure 1: EM signal processing performed in the reactive near-field (top) or in the radiative near-field (or far-field) (bottom) regions of the sources using a reconfigurable scattering device.
Let $\cal{S}$} denote the space occupied by the reconfigurable scattering device, whose configuration can be modified through a set $\bm{\theta}$ of parameters according to the specific implementation technology. Typically, the scattering device is a linear and passive structure composed of a large number of elementary reconfigurable scattering elements, enabling wave processing with significantly reduced power consumption. As will be illustrated in Section V, an antenna connected to a reconfigurable load provides a simple example of a reconfigurable scattering element. The purpose of the reconfigurable scattering device is to manipulate the EM field generated in the source space to produce a desired EM field in the receiver space. The processing occurs over the air, as the received signal results from the coupling among the scattering elements and their interaction with the sources themselves.
The scenario shown in Fig. 1 can encompass various ESP technologies and geometric configurations. Specifically, ESP solutions can generally be categorized into two types:
@@
  • Reactive region configuration: The scattering device is located in the reactive region of the source space (typically limited to a few wavelengths) so that inductive coupling takes place between them (see Fig. 1, top) and it can support the implementation of signal processing in the EM domain. Its role includes enabling large antenna arrays while minimizing the number of RF chains, reducing power consumption, and lowering latency compared to traditional MIMO architectures. The dual scheme where the scattering device is located close to the receiver space is also possible, thus moving the EM signal processing to the receiver side.
@@
  • Radiative near and far field region configuration: The reconfigurable scattering device is positioned in the radiative region of the source and receiver spaces and it is used to control the propagation environment (see Fig. 1, bottom). This approach offers a cost-effective alternative to deploying additional access points. In some configurations, the source and receiver spaces might match (see Section V-C).
Regardless of the specific category or underlying technology, the core ESP problem can be formulated as follows: Given a set of sources located in the source space $\mathcal{S}_{\text{T}}$, determine the configuration $\bm{\theta}$ of the scattering device that modifies the generated wave to obtain the desired EM field at the receiver space $\mathcal{S}_{\text{R}}$, while satisfying the constraints imposed by EM laws and the selected technology. For instance, one might require that the structure performs a multifunctional processing such that the electrical field ${\mathsf{E}}({\mathbf{r}})$, observed in the receiver space $\mathcal{S}_{\text{R}}$, is a linear combination of different arbitrary spatial responses $f_{n}\left({\mathbf{r}}\right)$, with ${\mathbf{r}}\in\mathcal{S}_{\text{R}}$, associated with the information $x_{n}$ transmitted by the $n$-th source. Denote by $\tilde{f}_{n}\left(\mathbf{r};\bm{\theta}\right)$ the actual response of the reconfigurable scattering device associated with the $n$-th source obtained with the configuration parameters $\bm{\theta}$. The problem is to configure the scattering device such that the actual responses are a good approximation of the desired ones according to some fidelity measure, and hence
${\mathsf{E}}({\mathbf{r}})=\sum_{n=1}^{N_{\mathrm{A}}}\tilde{f}_{n}\left(\mathbf{r};\bm{\theta}\right)\,x_{n}\simeq\sum_{n=1}^{N_{\mathrm{A}}}f_{n}\left(\mathbf{r}\right)\,x_{n}\,$, (1)
where $N_{\mathrm{A}}$ denotes the number of active sources. In other words, the challenge is to obtain different spatial responses simultaneously using the same configuration $\bm{\theta}$.
A further step involves making the set of parameters $\bm{\theta}$ dependent on data generated internally by the scattering device itself (e.g., when it incorporates a sensor in an Internet of things (IoT) application). This allows the information to be directly embedded into the scattered signal, eliminating the need for additional RF signal generation. As a result, this approach can lead to substantial energy savings on the sensor side [16, 17, 18].
@@
Notably, the concept of reconfigurable scatterers has a long and well-established history in the field of antenna theory. A pivotal milestone and precursor to modern solutions is R. Harrington’s theory of loaded scatterers [19], which is closely linked with characteristic mode decomposition. This theory involves configuring scatterers such that the principal radiating mode of the EM structure, when illuminated by the source wave, aligns with the desired response. Harrington later extended this theory to reactively controlled arrays, which represent one of the earliest realizations of reconfigurable scattering systems and correspond to the first category illustrated in Fig. 1 (top). In this context, passive scattering elements are positioned near a radiating antenna to exploit mutual coupling within the reactive near-field region, thereby modifying the antenna’s radiation characteristics [20].
@@
In recent years, these concepts have been revisited and advanced thanks to technological and material innovations (e.g., metamaterials), enabling the development of new EM structures that address the increasing demand for enhanced performance, flexibility, and cost efficiency in future wireless systems. These include XL-MIMO arrays (also known as extremely large aperture arrays) [21], large intelligent surfaces or holographic multiple-input multiple-output (HMIMO) surfaces [22, 23], dynamic metasurface antennas [6], fluid and reconfigurable antennas [24], and the latest developments in reactively controlled arrays, such as electrically steerable passive array radiators and dynamic scattering arrays [25, 26]. Significant advancements have concurrently been made in the optical community with the development of all-optical diffractive deep neural networks [27], and a similar concept has recently been applied to RF systems [28], inspiring the development of the stacked intelligent metasurface (SIM) technology, even though this technology is currently limited to linear processing capabilities [29]. Additionally, the introduction of novel volumetric metamaterials has enabled three-dimensional (3D) EM processing functions, such as first-derivative computation [30].
A notable example of new metamaterials is given by metasurfaces, which are bi-dimensional (2D) structures, either homogeneous or inhomogeneous, composed of interacting sub-wavelength-sized cells called meta-atoms. These structures may provide control of the reflected and transmitted EM field [31, 32]. Both metasurfaces and more conventional solutions, such as backscattering antenna arrays, have facilitated the development of reconfigurable intelligent surfaces, which have garnered significant interest within the scientific community. RISs are primarily used to manipulate the propagation environment and are considered a cornerstone of the emerging smart radio environments paradigm, whose aim is to integrate the propagation environment into the system’s design and optimization loop [33]. Specifically, RISs belong to the second category, shown in Fig. 1 (bottom), where they are typically employed to control the propagation environment [33, 34].
To build a general framework for the description of ESP, the foundational EM principles, on which the subsequent developments are based, are introduced in Section III. To this end, the preliminaries on the Green’s operator and the wavenumber representation are introduced in Section III-A, and then leveraged in Section III-B to define the concept of communication modes, first between a transmitting and a receiving space, and then in the presence of a programmable EM scatterer. This progression enables the definition of the mode transfer matrix for the scattering devices (Section III-C), ultimately culminating in the formal expression of the ESP problem. To account for realistic constraints on the types of processing that can be implemented at the EM level, Section IV analyzes the fundamental performance limits, building on the EM framework developed in Section III. Specifically, the wavenumber representation and essential signal theory concepts are used to derive the degrees-of-freedom (DoF) in both unbounded and bounded spaces (i.e., considering spatial constraints related to the transmitter/receiver/scatterer components), thereby linking the ESP problem previously formulated to its physical feasibility and fundamental limitations. The connection between electromagnetic and signal processing theory is provided through the circuit-based equivalent representation in Section V.A. Circuit theory represents one approach to building a physically consistent model of the system on which ESP algorithms can be designed. Finally, concrete examples of enabling EM technologies and over-the-air signal processing methodologies are provided in the rest of Section V. Some of the processing functions that can be achieved include over-the-air computation of MIMO EM precoding, discrete Fourier transform (DFT) computation, and direction-of-arrival (DoA) signal estimation.